# On the foundations and applications of Lorentz-Finsler Geometry

### v1

@@ 2.4.5. Anisotropic connections of Berwald and Chern
A semi-Finsler metric naturally yields two anisotropic connections:
@@
  • Berwald anisotropic connection: its Christoffel symbols are $\mathring{\Gamma}_{ij}^{a}(x,y)$ as in (11) and they can be defined for any nonlinear connection (in our case, they came from the formal ones $\gamma_{ij}^{k}(x,y)$ obtained from the fundamental tensor $g_{ij}(x,y)$ formally as in semi-Riemannian geometry, which appeared in the geodesic spray (7), thus in (8), (10)).
@@
    When the anisotropic Berwald connection is affine, the (semi-) Finsler manifold is called Berwald. In the positive definite case, Szabó [159] showed that the associated affine connection of a Berwald manifold is the Levi-Civita one of a Riemannian metric (see [66] for the indefinite case).
@@
  • Chern anisotropic connection: its Christoffel symbols are computed as the formal ones [^12] Notice that the $\gamma_{ij}^{k}(x,y)$’s do not satisfy the cocycle rule in (14), but the transformation obtained replacing in the left hand side therein $\bar{y}(x,y)$ by $y$. $\gamma_{ij}^{k}(x,y)$ but now replacing the derivatives $\partial/\partial{x^{i}}$ by the corresponding derivatives $\delta/\delta x^{i}$ in (9). It plays a role similar to the Levi-Civita connection in semi-Riemannian geometry, as it is the unique symmetric anisotropic connection parallelizing $g$, [90, Sect. 4.1]. Another interpretation in terms of $\nabla^{Z}$ can be found in [91, Proposition 3.9]; in particular, if $Z$ is taken parallel at a point $x$ (a property which depends only on the nonlinear connection, see (12)), then the Christoffel symbols of the semi-Riemannian metric $g_{Z}$ agree at $x$ with those of Chern’s anisotropic connection at $(x,Z_{x})$, see [154, formula (7.17)].
@@
    The difference between the Chern and Berwald anisotropic connections is (up to metric equivalence) the so-called Landsberg tensor. When this tensor vanishes, the semi-Finsler manifold is called Landsberg.
@@
Explicitly, the Berwald and Landsberg tensors, as well as the contraction of the latter called mean Landsberg, are written in coordinates, starting at the geodesc spray in (7), respectively as follows:
$B^{\mu}_{\nu\rho\sigma}:={G}^{\mu}_{\cdot\nu\cdot\rho\cdot\sigma},\qquad{\mathrm{Lan}_{\mu\nu\rho}}:=\frac{1}{2}y_{\alpha}B^{\alpha}_{\mu\nu\rho},\qquad{\mathrm{Lan}_{\mu}}:=g^{\alpha\beta}\mathrm{Lan}_{\alpha\beta\mu}$ (15)
@@
where the dot symbol denotes vertical derivative (that is, $\cdot\alpha$ means $\partial_{y^{\alpha}}$). Notice that $y_{\alpha}:=y^{\beta}g_{\alpha\beta}$, showing the Landsberg dependence on the semi-Finsler metric and not only on the geodesic spray.
@@
Summarizing, these tensors vanish for semi-Riemannian manifolds, the Berwald tensor measures at what extent the nonlinear connection is truly non-linear (i.e., non-affine) and the Landsberg one at what extent the anisotropic Chern connection is not determined only by the non-linear connection.
@@ 2.4.6. Finsler connections
    Chern: $H_{ij}^{k}=\Gamma_{ij}^{k}$ (Christoffel symbols for Chern), $V_{ij}^{k}=0$.
@@
    Cartan: $H_{ij}^{k}=\Gamma_{ij}^{k}$, $V_{ij}^{k}=C_{ijk}^{a}$.
Berwald and Chern linear connections are vertically trivial (i.e. $V_{ij}^{k}=0$), thus, equivalent to their anisotropic counterparts.
@@ 2.4.7. Curvatures
$\mathcal{R}_{ij}^{k}=\delta_{j}N_{i}^{k}-\delta_{i}N_{j}^{k},\qquad\mathrm{Ric}=y^{b}\,\mathcal{R}_{ba}^{a},\qquad\mathrm{Tor}_{ij}^{k}=\partial_{y^{j}}N_{i}^{k}-\partial_{y^{i}}N_{j}^{k}.$ (16)
The tensor $\mathcal{R}_{\;j}^{k}=y^{b}\mathcal{R}_{bj}^{k}$, whose contraction yields the Ricci scalar, may be regarded as a predecessor of the flag curvature below (see [8, Ex. 2.5.7]). By (8), the torsion vanishes for the nonlinear connection from a geodesic spray.
@@
Note that $K(x,y,\Pi)$ depends on the direction (flagpole) $y$ and the plane $\Pi=\mathrm{span}\{y,v\}$ (flag) but not on the choice of $v$. The Ricci scalar $\mathrm{Ric}(x,y)$ can be computed as an average of the flag curvatures with flag pole $y$. In the literature, a Ricci tensor $\mathrm{Ric}_{ij}(x,y)$ introduced by Akbar-Zadeh and containing the same information of the Ricci scalar is also used [8, p. 192]. Moreover, a weighted Ricci curvature was introduced by Ohta [138, 140], in order to adapt the well-known N-Bakry-Émery-Ricci tensor on a weighted Riemannian manifold (see for example [112]) to the Finsler setting. It plays a relevant role in comparison theorems and yields splitting results [139] extensible to the Lorentz-Finsler case [116].
@@ 3.1. Finsler globally hyperbolic splittings
In this subsection, the main aim is to achieve the Finslerian splittings in Theorems 3.1, 3.3 ensuring also additional properties listed below (Remark 3.4). Reference [150] widely expands our Lorentzian setting here and other Finsler developments.
@@ 3.1.1. The Lorentzian splitting
Classical Geroch’s theorem [68] asserts the equivalence, in the Lorentz case, between the global hyperbolicity of a spacetime $(M,g)$ and the existence of a topological acausal Cauchy hypersurface $S$. What is more, he constructed a Cauchy time function $t$ with all the levels $t=$ constant Cauchy hypersurfaces (as $S$) and, then, a homeomorphism between $M$ and $\mathds{R}\times S$. The long-standing question whether $S$ or even the Cauchy time $t$ could be found smooth, was answered in the positive by Bernal and the author in [24] and [25], respectively. The last reference, however, not only proved the existence of a smooth Cauchy $t$ but also the existence of a Cauchy temporal function $\tau$ (i.e, smooth with timelike gradient $\nabla\tau$). This not only implies the improvement of the topological splitting $\mathds{R}\times S$ into a smooth one with spacelike Cauchy hypersurfaces, but also the existence of an associated smooth global orthogonal splitting, that is, with no cross terms between the $\mathds{R}$ and $S$ parts. On the one hand such splitting admits first-principles implications about the possibility to recover globally the notions of time and space [152] and, on the other, Cauchy temporal functions $\tau$ also exhibit more technical advantages:
@@
    As proven by Burstcher and García-Heveling [35], $\tau$ can be chosen $h$-steep with respect to a complete Riemannian metric (a notion introduced by Bernard and Suhr [28] see below), which is a condition more restrictive than being Cauchy temporal; additionally, it can be found steep, see [34, Appendix C]).
    This is important in the context of Sormani and Vega’s null distance $\hat{d}$ associated with $\tau$, see [156]. Indeed, $h$-steepeness implies the completeness of $\hat{d}$, a crucial property for the theory of convergence of metrics. Moreover, $h$-steepness also becomes equivalent to the completeness of the Wick rotated metric of the conformal metric $g/|\nabla\tau|^{2}$, which also has implications for this theory [34, Sect. 4].
@@
    In case of analytic metrics, the Cauchy temporal function as well as the associated splitting, can be chosen also analytic, see [150, Th. 2.10]. Obviously, the extensibility from a compact hypersurface $S_{0}$ to a Cauchy hypersurface seen in item seen in item (2) cannot hold. However, the smooth extension $S$ obtained in that item can be approximated by anaylitic Cauchy spacelike hypersurfaces.

### v2

@@ 2.4.5. Anisotropic connections of Berwald and Chern
A semi-Finsler metric naturally yields two anisotropic connections defined as in (15), in terms of Christoffel symbols (satisfying (14)):
@@
  • Berwald anisotropic connection. ts Christoffel symbols are $\mathring{\Gamma}_{ij}^{a}(x,y)$ as in (12) and they can be defined for any nonlinear connection $N^{a}_{i}$. In our case, recall that $N^{a}_{i}$ came from the formal ones $\gamma_{ij}^{k}(x,y)$ obtained from the fundamental tensor $g_{ij}(x,y)$ as in semi-Riemannian geometry (indeed, they appeared in the spray (8), thus in (9), (11)).
@@
    When the anisotropic Berwald connection is affine, the (semi-) Finsler manifold is called Berwald. This is characterized by the vanishing of the Berwald tensor (positive homogeneous of degree -1), obtained by taking the vertical derivative of the Berwald anisotropic connection. In the positive definite case, Szabó [174] showed that the associated affine connection of a Berwald manifold is the Levi-Civita one of a Riemannian metric (see [69] for the indefinite case).
@@
  • Chern anisotropic connection. its Christoffel symbols $^{C}\Gamma_{ij}^{a}$ are computed as the formal ones [^10] Notice that the $\gamma_{ij}^{a}(x,y)$’s do not satisfy the cocycle rule in (14), but the transformation obtained replacing in the left hand side therein $\bar{y}(x,y)$ by $y$. $\gamma_{ij}^{a}(x,y)$ but now replacing the derivatives $\partial/\partial{x^{i}}$ by the corresponding derivatives $\delta/\delta x^{i}$ in (10). It plays a role similar to the Levi-Civita connection in semi-Riemannian geometry. Indeed, it is the unique symmetric anisotropic connection parallelizing $g$, [98, Sect. 4.1]. Another interpretation in terms of $\nabla^{Z}$ can be found in [99, Proposition 3.9]; in particular, if $Z$ is taken parallel at a point $x$ (a property which depends only on the nonlinear connection, see (13)), then the Christoffel symbols of the semi-Riemannian metric $g_{Z}$ agree at $x$ with those of Chern’s anisotropic connection at $(x,Z_{x})$, see [168, formula (7.17)].
@@
    The difference between the Chern and Berwald anisotropic connections is (up to metric equivalence) the so-called Landsberg tensor. When this tensor vanishes, the semi-Finsler manifold is called Landsberg. (The vertical derivative of the Chern anisotropic connection also yields a tensor, which is the sum of the Berwald tensor and the vertical derivative of the Landsberg one.)
@@
Explicitly, the Berwald and Landsberg tensors, as well as the contraction of the latter called mean Landsberg, are written in coordinates (starting at the geodesic spray in (8)), respectively as follows:
[…]
@@
where the dot symbol denotes vertical derivative (that is, $\cdot\alpha$ means $\partial_{y^{\alpha}}$). Notice that $y_{a}:=y^{b}g_{ab}$, showing the Landsberg dependence on the semi-Finsler metric and not only on the geodesic spray. Moreover, the difference between the anisotropic connections $\mathring{\Gamma}_{ij}^{a}$ and $^{C}{\Gamma}_{ij}^{a}$ becomes $g^{ab}\mathrm{Lan}_{ijb}$.
@@
Summarizing, these tensors vanish for semi-Riemannian manifolds, the Berwald tensor measures to what extent the nonlinear connection is truly non-linear (i.e., non-affine) and the Landsberg one to what extent the anisotropic Chern connection is not determined only by the non-linear connection.
@@ 2.4.6. Finsler connections
    Chern: $H_{ij}^{k}=^{C}\Gamma_{ij}^{k}$ (Christoffel symbols for Chern), $V_{ij}^{k}=0$.
@@
    Cartan: $H_{ij}^{k}=^{C}\Gamma_{ij}^{k}$, $V_{ij}^{k}=C_{ijk}^{a}$.
Berwald and Chern linear connections are vertically trivial (i.e. $V_{ij}^{k}=0$), thus, equivalent to their anisotropic counterparts. Indeed, the (anisotropic) Christoffel symbols $\Gamma_{ij}^{k}(x,y)$ for the latter ones can be regarded as the corresponding (linear) Christoffel symbols for the former ones (see formula (30) in [165], where $\Delta$ therein vanishes for a vertically trivial connection). Given a curve $s\mapsto Z(s)=(x(s),y(s))$ in $A$, one has naturally a parallel transport associated with the linear connection, as well as an anisotropic parallel transport associated with $\nabla^{Z}$, the latter with further possibilities see [110, Sect. 7].
@@ 2.4.7. Curvatures
$\mathcal{R}_{ij}^{a}=\delta_{j}N_{i}^{a}-\delta_{i}N_{j}^{a},\qquad\mathrm{Ric}=y^{b}\,\mathcal{R}_{ba}^{a},\qquad\mathrm{Tor}_{ij}^{a}=\partial_{y^{j}}N_{i}^{a}-\partial_{y^{i}}N_{j}^{a}.$ (18)
The tensor $\mathcal{R}_{\;j}^{k}=y^{b}\mathcal{R}_{bj}^{k}$, whose contraction yields the Ricci scalar, may be regarded as a predecessor of the flag curvature below (see [10, Ex. 2.5.7]). By (9), the torsion vanishes for the nonlinear connection from a geodesic spray. It is worth noticing
$\left[\frac{\delta}{\delta x^{i}},\frac{\delta}{\delta x^{j}}\right]=\;\mathcal{R}_{ij}^{a}\;\frac{\partial}{\partial y^{a}}\;,$
so that the horizontal distribution is integrable around a point $(x,y)\in\bar{A}$ if and only if $\mathcal{R}_{ij}^{a}\equiv 0$ around $(x,y)$.
@@
Note that $K(x,y,\Pi)$ depends on the direction (flagpole) $y$ and the plane $\Pi=\mathrm{span}\{y,v\}$ (flag) but not on the choice of $v$. The Ricci scalar $\mathrm{Ric}(x,y)$ can be computed as an average of the flag curvatures with flag pole $y$. In the literature, a Ricci tensor $\mathrm{Ric}_{ij}(x,y)$ introduced by Akbar-Zadeh and containing the same information of the Ricci scalar is also used [10, p. 192]. Moreover, a weighted Ricci curvature was introduced by Ohta [147, 149], in order to adapt the well-known N-Bakry-Émery-Ricci tensor on a weighted Riemannian manifold (see for example [120]) to the Finsler setting. It plays a relevant role in comparison theorems and yields splitting results [148] extensible to the Lorentz-Finsler case [125].
2.5. Symplectic and contact viewpoints
This part follows the classical setting in Mechanics [2, 6], adapted when necessary to the Finsler [52] and Lorentz-Finsler [85] cases. Notice that the usual notation in Mechanics for coordinates in the bundles $TM$ and $T^{*}M$ are rewriten here as $q^{i},\dot{q}^{i}\equiv x^{i},y^{i}$ and $q^{i},p_{i}\equiv x^{i},p_{i}$, resp. Morever, in Finslerian Mechanics one tipically uses $TM\setminus\mathbf{0}$ and its dual bundle and we will be particularly interested in the set of causal vectors $\bar{A}\subset TM\setminus\mathbf{0}$ and its dual $\bar{A}^{*}$ of causal forms.
2.5.1. Basic Hamiltonian and Lagrangian approach
The cotangent bundle $\hat{\pi}:T^{*}M\rightarrow M$ admits the tautological form, written in bundle coordinates $(x^{i},p_{i})$ (with sum in repeated indexes)
$\theta=p_{i}dx^{i},\quad\hbox{that is},\quad\theta(y^{i}\partial_{x^{i}}+\dot{p}_{j}\partial_{p_{j}})|_{(x,p)}=p(y^{i}\partial_{x^{i}}|_{x})=p_{i}y^{i},$
where $(x^{i},p_{i},y^{i},\dot{p}_{i})$ are coordinates in $T(T^{*}M)$ and $y^{i}\partial_{x^{i}}=\hat{\pi}_{*}(y^{i}\partial_{x^{i}}+\dot{p}_{j}\partial_{p_{j}})$, making $\theta$ independent of coordinates. Then, the Poincaré 2-form,
$\omega=-d\theta=dx^{i}\wedge dp_{i}.$
is symplectic i.e., closed with $\omega\wedge\dots^{(n)}\wedge\omega\neq 0$. By Darboux theorem, every symplectic form looks like as above in suitable coordinates. The fiber of $TM^{*}$ at each $p\in M$ is a Lagrangian submanifold, i.e., $\omega\equiv 0$ on the submanifold and it has maximum dimension $n$.
Moreover, $\omega$ also yields a linear isomorphism (symplectic gradient)
$T(T^{*}M)\longrightarrow T^{*}(T^{*}M),\quad X\mapsto i_{X}\omega(\equiv\omega(X,\cdot)).$
Notice that this map restricts naturally when $T^{*}M$ is replaced by $\bar{A}^{*}$. Thus, it permits to define, for any function $H:\bar{A}^{*}\subset T^{*}M\rightarrow\mathds{R}$ its Hamiltonian vector field $X_{H}$, with integral curves $t\to(x(t),p(t))$, characterized by
$\begin{array}[]{c}dH=i_{X_{H}}\omega,\;\;\;\hbox{then,}\;\;\;X_{H}=\frac{\partial H}{\partial p_{i}}\frac{\partial}{\partial x^{i}}-\frac{\partial H}{\partial x^{i}}\frac{\partial}{\partial p_{i}},\;\hbox{and }\;y^{i}=\frac{\partial H}{\partial x^{i}},\,\dot{p}_{i}=-\frac{\partial H}{\partial x^{i}},\end{array}$
the last pair being the Hamilton equations. The flow of $X_{H}$ preserves the symplectic structure and the Hamiltonian $H$ becomes constant along the integral curves of $X_{H}$.
A (2-homogeneous) Finsler or Lorentz-Finsler metric must be regarded as a Lagrangian $L/2:\bar{A}\subset TM\rightarrow\mathds{R}$ (the factor 1/2 being convenient). It permits to define the generalized momenta in coordinates
$p_{i}:=\frac{1}{2}\frac{\partial L}{\partial y^{i}}=\frac{1}{2}\frac{\partial g_{jk}}{\partial y^{i}}y^{j}y^{k}+g_{ik}y^{k},$
so that $p_{i}dx^{i}$ recovers the Legendre map (Def. 2.10) in our case:
$\flat:\bar{A}\rightarrow\bar{A}^{*},\qquad y^{j}\partial_{x^{j}}|_{x}\mapsto p_{i}dx^{i}|_{x}=g_{ij}(x,y)y^{j}dx^{i}|_{x},$ (19)
just noticing that, by the 0-homogeneity of $g_{ij}$ and Euler’s theorem,
$\frac{1}{2}\frac{\partial g_{jk}}{\partial y^{i}}y^{j}y^{k}\left(dx^{i}(y^{l}\partial_{x^{l}})\right)=\frac{1}{2}y^{j}y^{k}\left(\frac{\partial g_{jk}}{\partial y^{i}}y^{i}\right)=0.$
The standard Hamiltonian $H_{L}$ associated with $L$ is its Legendre transform,
$H_{L}(x,p)=p_{i}\,y^{i}(x,p)-\frac{1}{2}L(x,y(x,p))=\frac{1}{2}L(x,y(x,p)).$
Thus $H_{L}=L\circ\sharp$ and the Legendre map sends bijectively the solutions $\gamma$ of the Lagrange equations (5) into solutions $(\gamma^{\prime})^{\flat}$ of Hamilton’s preserving the (kinetic and total) energy $E=L(\gamma^{\prime})/2=H((\gamma^{\prime})^{\flat})$.
2.5.2. The Hilbert form
Formula (19) can be alternatively interpreted as a 1-form $\eta_{g}$ on $\bar{A}$, the Hilbert form, defined as the pullback [^11] In the Finsler case some times this is divided by $F(x,y)$. However, this would introduce an issue in the Lorentz-Finsler case for lightlike geodesics.
$\eta_{g}:=\flat^{*}\theta=g_{ij}(x,y)y^{i}dx^{j}|_{(x,y)}.$
In a natural way, $\omega_{g}=-d\eta_{g}$ provides a symplectic form on $T\bar{A}$ (see [52, Prop.7.13]). Then, the $\omega_{g}$-Hamilton equations for $L$ (regarded as a Hamiltonian function) are equivalent to the geodesic equations for $(M,L)$, which turns out the Lagrange equations for the Lagrangian $L/2$. Indeed, the $L$-geodesic vector field $G$ on $\bar{A}$ in (8) is the $\omega_{g}$-Hamiltonian one $X_{L}$, see [52, Proposition 7.19].
As the integral curves of $G$ are the velocities of the geodesics, $\eta_{g}(G|_{(x,y)})=g_{ij}(x,y)y^{i}y^{j}$ is constant along each geodesic and vanishes only on the lightlike ones. In particular, $G$ is tangent to the indicatrix $\Sigma=L^{-1}(1)$ and $\eta_{g}(G)|_{\Sigma}\equiv 1$ for both the Finsler and Lorentz-Finsler cases.
2.5.3. Contact geometry
$\Sigma$ is a hypersurface of $TA$ with odd dimension 2n12n-1, where the standard contact setting for the Finsler case [52] extends to the Lorentz-Finsler one [85].
A contact form $\eta$ is a 1-form satisfying $\eta\wedge d\eta\wedge\dots\wedge^{(n-1)}d\eta\neq 0$ so that ker $\eta$ is a contact structure, that is, a distribution of hyperplanes which is maximally integrable, in the sense that the distribution can be locally expressed from some $\eta$ as above. We will be interested in the case that this expression is global (i.e., the distribution is co-orientable). Anyway, the contact structure is also determined by $f\eta$, with $f\neq 0$ (a non-vanishing function). The primary example of contact form is the spherical cotangent bundle $ST^{*}M$ of $M$ (i.e. the quotient of $T^{*}M\setminus\{0\}$ by the $\mathds{R}_{>0}$ action of positive homotheties), where the tautological 1-form $\theta$ induces the required hyperplane distribution. The fibers of this bundle are Legendrian submanifolds, i.e., the contact form vanish on them and they have the maximum dimension n1n-1.
A contact form provides a unique Reeb vector field, characterized by
$\eta(R)=1,\qquad i_{R}d\eta=0,$
and, moreover, for each function $f$ on $\Sigma$, the vector field $X_{f}$ satisfying:
$\eta(X_{f})=f,\qquad i_{X_{f}}d\eta=df(X_{f})\alpha-df.$
$X_{f}$ is a contact vector field, i.e., its flow is composed by contactomorphisms (which preserve the contact structure but not necessarily the contact form). Notice that, when $f\not=0$, then $X_{f}$ is the Reeb vector of $\alpha/f$.
The Hilbert form $\eta_{g}$ vanishes on the the radial vector field $y^{i}\partial_{y^{i}}$, which is transversal to $\Sigma$. Its restriction $\eta$ to $\Sigma$ has rank 2n22n-2 and becomes the natural contact form $\eta$ on $\Sigma$ (see [85, Example 2.1.1] and references therein). Its Reeb vector $R$ is just the restriction of the geodesic field $G$ to $\Sigma$. Moreover, one can check that $\eta=R^{\flat}$, where the isomorphism $\flat:TA\to TA^{*}$ is computed using the Sasaki metric (see [52, Sect. 4.1]) restricted to $\Sigma$.
Notice that $\eta_{g}$ can also be restricted to $\mathcal{C}$, however, the setting changes as $\eta_{g}(G)|_{\mathcal{C}}\equiv 0$). We postpone its study to the space of cone geodesics in §3.2.
2.5.4. About the Hamilton-Jacobi equation
When studying the propagation of light in a classical setting, the relation between particle trajectories and wavefronts is well established in classical Mechanics from Hamilton-Jacobi equation (see [2, 6] or the thesis [178]).
When considering light propagation or shortest paths in Riemannian geometry and geometric optics, Hamilton-Jacobi reduces to the eikonal equation, namely, $|\nabla u|(x)=n(x)$, where $n(x)$ is a positive function, tipycally either $n(x)\equiv 1$ or the refraction index $n(x)=c/v(x)$ —the quotient between the speed of light in vacuum and in (isotropic) media.
The function $n(x)$ can be regarded as a cost function to be locally minimized by the trajectories, so that the wavefronts become the level sets of an appropriate distance function. This setting will underlie in the applications to both Lorentz Geometry and Classical Mechanics to be developed next.
@@ 3.1. Finsler globally hyperbolic splittings
In this subsection, the main aim is to achieve the Finslerian splittings in Theorems 3.1, 3.3 ensuring also additional properties listed below (Remark 3.4). Reference [163] expands our Lorentzian setting here and other Finsler developments.
@@ 3.1.1. The Lorentzian splitting
Classical Geroch’s theorem [72] asserts the equivalence, in the Lorentz case, between the global hyperbolicity of a spacetime $(M,g)$ and the existence of a topological acausal Cauchy hypersurface $S$. What is more, he constructed a Cauchy time function $t$ with all the levels $t=$ constant Cauchy hypersurfaces (with the same properties as $S$) and, then, a homeomorphism between $M$ and $\mathds{R}\times S$. The long-standing question whether $S$ or even the Cauchy time $t$ could be found smooth, was answered in the positive by Bernal and the author in [26] and [27], respectively. The last reference, however, not only proved the existence of a smooth Cauchy $t$ but also the existence of a Cauchy temporal function $\tau$ (i.e, smooth with timelike gradient $\nabla\tau$). This not only implies the improvement of the topological splitting $\mathds{R}\times S$ into a smooth one with spacelike Cauchy hypersurfaces, but also the existence of an associated smooth global orthogonal splitting, that is, with no cross terms between the $\mathds{R}$ and $S$ parts. On the one hand such splitting admits first-principles implications about the possibility to recover globally the notions of time and space [164] and, on the other, Cauchy temporal functions $\tau$ also exhibit more technical advantages:
@@
    As proven by Burstcher and García-Heveling [37], the temporal $\tau$ can be chosen $h$-steep with respect to a complete Riemannian metric (a notion introduced by Bernard and Suhr [30] see below), which is a condition more restrictive than being Cauchy temporal (additionally, it can be found also steep, see [36, Appendix C]).
    This is important in the context of Sormani and Vega’s null distance $\hat{d}$ associated with $\tau$, see [171]. Indeed, $h$-steepeness implies the completeness of $\hat{d}$, a crucial property for the theory of convergence of metrics. Moreover, $h$-steepness also becomes equivalent to the completeness of the Wick rotated metric of the conformal metric $g/|\nabla\tau|^{2}$, which also has implications for this theory [36, Sect. 4].
@@
    In case of analytic metrics, the Cauchy temporal function, as well as the associated splitting, can be chosen also analytic, see [163, Th. 2.10]. Obviously, the extensibility from a compact hypersurface $S_{0}$ to a Cauchy hypersurface seen in item (2) cannot hold in an analytic way. However, the smooth extension $S$ obtained in that item can be approximated by analytic Cauchy spacelike hypersurfaces.